OpenStudy (anonymous):

find the area bounded by parabola y2=4x and y=2x-4

OpenStudy (anonymous):

i need detail explaination for this problem

OpenStudy (blockcolder):

Hint: \[\cos A \cos B=\frac{1}{2}(\cos(A-B)+\cos(A+B))\]

OpenStudy (anonymous):